OpenStudy (anonymous):

sinx. secx/1+x. tanx

OpenStudy (anonymous):

what is the equation of sec?

OpenStudy (anonymous):

so what is the actual equation above?

OpenStudy (anonymous):

sinx. secx/1+x. tanx, is the question

OpenStudy (anonymous):

Can u write it using equation tab at the bottom left of commenting box?

OpenStudy (anonymous):

ok i'll try

OpenStudy (anonymous):

sin x. sec x div 1+x. tan x

OpenStudy (anonymous):

is there anyone who can answer this

OpenStudy (ajprincess):

\[\frac{ \sin x.\sec x }{ (1+x)\tan x }\]

OpenStudy (ajprincess):

is that ur question?

OpenStudy (anonymous):

\[f(x) = \sin x. 1/\cos x \div 1+x. \sin x/\cos \]

OpenStudy (anonymous):

yes you got that

OpenStudy (ajprincess):

so what do u want to do? u wanna simplify?@KingKhan007

OpenStudy (ajprincess):

@KingKhan007

OpenStudy (anonymous):

I mean simplify this

OpenStudy (ajprincess):

\[\frac{\ \sin x\sec x }{ (1+x)\tan x }=\large\frac{ \sin x \frac{ 1 }{ cosx }}{ (1+x)\frac{\ \sin x }{ \cos x } }\] \[=\frac{ \sin x\cos x }{ (1+x)\sin x\cos x }\] cancel out the common factors

OpenStudy (anonymous):

sorry i didn't got that sinx and cosx are in division with each other

OpenStudy (anonymous):

can you please complete this for me

OpenStudy (ajprincess):

i meant to say giving out answers is not allowed.

OpenStudy (anonymous):

ok don't answer me but please simplify it for me little further

OpenStudy (ajprincess):

 Does this help? if u dnt understand pls do ask I will try my best to help u:) well giving out answers is not allowed in os. pls check the code of conduct for more info. this is the link for it. http://openstudy.com/code-of-conduct